How long would it take a ball dropped from an 81-foot tower to hit the ground? Round your answer to the nearest two decimal places.
Answer by scott8148(6628)

the ball is accelerated from rest by gravity

s = (1/2)at^2

sqrt[(2s) / a] = t

sqrt[(2 * 81) / 32] = t